A 1000 L tank containing 100 moles of compressed air at 235.4kPa of pressure at 5.40oC is heated by the sun to 24.0 oC. What is the resulting pressure in kPa?


1
Expert's answer
2021-02-23T10:04:07-0500

Assuming the volume does not change, the new pressure can be found from combined gas law:


"\frac{VP}{T}=\frac{VP_f}{T_f},\\\space\\
\frac{P}{T}=\frac{P_f}{T_f},\\\space\\
P_f=\frac{T_f}{T}P=\frac{273+24}{273+5.4}·235.4=521.1\text{ kPa}."